# Counterfactual Fairness through Transforming Data Orthogonal to Bias

### v1

@@ Abstract.
Machine learning models have shown exceptional prowess in solving complex issues across various domains. Nonetheless, these models can sometimes exhibit biased decision-making, leading to disparities in treatment across different groups. Despite the extensive research on fairness, the nuanced effects of multivariate and continuous sensitive variables on decision-making outcomes remain insufficiently studied. We introduce a novel data pre-processing algorithm, Orthogonal to Bias (OB), designed to remove the influence of a group of continuous sensitive variables, thereby facilitating counterfactual fairness in machine learning applications. Our approach is grounded in the assumption of a jointly normal distribution within a structural causal model (SCM), proving that counterfactual fairness can be achieved by ensuring the data is uncorrelated with sensitive variables. The OB algorithm is model-agnostic, catering to a wide array of machine learning models and tasks, and includes a sparse variant to enhance numerical stability through regularization. Through empirical evaluation on simulated and real-world datasets—including the adult income and the COMPAS recidivism datasets—our methodology demonstrates its capacity to enable fairer outcomes without compromising accuracy.
Counterfactual fairness, Data pre-processing, Algorithmic Fairness
1. Introduction
In recent years, machine learning has emerged as a pivotal technology, driving advancements across a broad spectrum of real-world applications, from healthcare diagnostics (Feng et al., 2022), hiring decision-making systems (Cohen et al., 2020), to loan assessments (Mukerjee et al., 2002). Its ability to learn from and make predictions or decisions based on large data sets has been transformative in addressing complex problems. However, the broader application prospects of some machine learning techniques is hampered by the implicit biases ingrained in the data it learns from, leading to outcomes that systematically and unfairly disadvantage certain groups (Bolukbasi et al., 2016; Haoyu Chen and Ghosh, 2023; Kilbertus et al., 2020; Andrus et al., 2021). This issue of bias in machine learning models is not merely a technical challenge but a fundamental concern that threatens to reinforce existing social inequalities.
@@ 1. Introduction
Such fairness concern in machine learning has catalyzed a growing body of research aimed at identifying, understanding, and mitigating biases present in data and algorithms. Among the various conceptual frameworks developed to address this issue (Bolukbasi et al., 2016; Hardt et al., 2016; Dwork et al., 2012; Grgic-Hlaca et al., 2016; Zemel et al., 2013; Zafar et al., 2017), the notion of counterfactual fairness (Kusner et al., 2017) stands out. Counterfactual fairness seeks to ensure that a decision made by a machine learning model would remain unchanged if a sensitive (or protected) variable of an individual were different, all else being equal. This concept is particularly powerful as it aligns closely with intuitive notions of individual fairness and justice, offering a rigorous standard against which to measure and rectify bias.
Numerous previous research efforts have focused on attaining counterfactual fairness, making significant strides in this area (Kusner et al., 2017; Wang et al., 2019; Hardt et al., 2016; Haoyu Chen and Ghosh, 2023; Chiappa, 2019). A key example is the study by (Wang et al., 2019), which targets non-discrimination through adjusting the predictions based on the empirical joint distribution of the data, aiming to achieve equal bias and accuracy across different groups. However, these approaches typically presuppose a complete understanding of the causal relationships among all variables without unmeasured confounders and consider sensitive variables to be binary (or categorical), so that they can be easily isolated or adjusted. As a result, traditional methods of fair learning face challenges in addressing situations involving multivariate and continuous sensitive variable with intricate inter-dependency.
@@
Figure 1. This motivating example illustrates the distinction between binary and continuous sensitive variables in the context of hiring decisions. The dashed lines indicate the predicted hiring decisions and the shaded area indicate the unbiased true decisions. The left panel simplifies the scenario with a binary sensitive variable, such as Gender, where adjustments for fairness are more straightforward due to the clear dichotomy in data. The right panel, however, delves into the complexity introduced by a continuous sensitive variable, like socio-economic status, demonstrating the intricate task of achieving fairness across a spectrum.
Take the employment hiring process as an instance shown in Figure 1, socioeconomic status can’t be easily classified into a small number of discrete categories. It is a complex, multidimensional variable influenced by education, income, occupation, and even subtler factors like neighborhood and parental education. A candidate’s profile is a complex amalgamation of their experiences, skills, educational background, and personal characteristics. A model trained on historical hiring data might inadvertently learn to favor candidates from prestigious universities or with certain types of work experiences—criteria that are often correlated with socioeconomic status. This form of implicit bias emerges either because the data used for training lacks comprehensive representation of all possible candidate profiles or because traditional approaches to fairness do not adequately address the subtle interplay and impact of various sensitive variables on decision-making outcomes.
@@
To tackle these challenges, we develop a novel data pre-processing approach that aims to remove the influence of a group of continuous sensitive variables from the data, thereby ensuring counterfactual fairness in subsequent machine learning tasks. We first prove that the counterfactual fairness can be attainable easily by making the data uncorrelated with the group of sensitive variables, based on the assumption of a jointly normal distribution within a structural causal model (SCM) framework (Pearl, 2009). This assumption is applicable across a broad spectrum of applications where standardization of data is required, ensuring that all variables, both sensitive and non-sensitive, are normalized to have a mean of zero and a variance of $1$. Motivated by this understanding, we consider all sensitive variables collectively and propose a data pre-processing algorithm, referred to as Orthogonal to Bias (OB). This algorithm is designed for minimal data adjustments to achieve orthogonality between non-sensitive and sensitive data. To facilitate numerical stability, we also present a sparse variant of this algorithm which incorporates a regularization term. Then the resulting data is ready to serve as input for machine learning models in downstream tasks without being influenced by the undesirable bias associated with the complexities of sensitive variables. It is also important to note that our proposed algorithm is model-agnostic, making it suitable for a variety of machine learning models and tasks. Lastly, we evaluate our algorithm’s performance on a simulated data set and two real-world data sets, including the adult income data and the COMPAS recidivism data, demonstrating that our approach enables machine learning models to achieve fairer outcomes with comparable accuracy to current state-of-the-art fair learning methods. In the numerical results, we found that our approach is not limited by the assumptions about data distribution, indicating its applicability to a wider range of scenarios.
@@
    We show that achieving counterfactual fairness is feasible by ensuring orthogonality between non-sensitive and sensitive data when they are jointly normal.
  2. (2)
@@
    We introduce a model-agnostic data-pre-processing algorithm, termed as Orthogonal to Bias (OB), which facilitates counterfactual fairness across a broad spectrum of downstream machine learning applications.
  3. (3)
@@
    We validate the enhanced efficacy of our algorithm compared to the existing state-of-the-arts through evaluations on both synthetic and real-world data sets.
(a) The SCM and the typical fair learning
(b) Fair learning by transforming data orthogonal to bias
@@
Figure 2. Illustration of (a) the structural causal model (SCM) and a common fair learning strategies, as well as (b) the proposed data pre-processing algorithm Orthogonal to Bias (OB). The white nodes $A$, $Y$, and $\hat{Y}$ are the non-sensitive variables, the decision variable, and its prediction, respectively. The red nodes $B$ represent the sensitive variable. The $\hat{A}$ is the transformed data that is orthogonal to bias in $B$. The gray nodes represent exogenous variables.
2. Literature Review
This work is related to several streams of algorithmic fairness literature which we review in this section.
Fairness in Machine Learning
@@ 2. Literature Review
The pursuit of fair decision-making in machine learning has led to diverse approaches for defining and quantifying fairness. Researchers commonly adopt either observational or counterfactual approaches to formalize fairness. Observational methods typically characterize fairness through metrics derived from observed data and predicted outcomes (Kamiran and Calders, 2012; Luong et al., 2011; Haoyu Chen and Ghosh, 2023; Wei et al., 2024). Metrics such as individual fairness (IF) (Dwork et al., 2012), demographic parity or Group Fairness (Zemel et al., 2013; Khademi et al., 2019) and equalized odds (Hardt et al., 2016) fall under this category. The key idea for the observational fairness metric is viewing fairness as treating similar individuals or individuals belonging to the same groups similarly. For example, IF defines fairness as treating any two individuals who are similar with respect to a particular task similarly (Zemel et al., 2013).
In contrast, counterfactual approaches proposes a causal approach to defining fairness. These definitions assess fairness based on how predictions would change if sensitive attributes were altered (Kusner et al., 2017; Wang et al., 2019; Hardt et al., 2016; Haoyu Chen and Ghosh, 2023; Chiappa, 2019). With the help of the potential outcome concept, the measuring of fairness is no longer restricted to the observable quantities. For instance, the Equal Opportunity (EO) definition, akin to individual fairness, directly compares the actual and counterfactual decisions of the same individual, rather than relying on comparisons between the observation of two similar individuals (Wang et al., 2019).
While observational definitions of fairness can be incorporated into optimization problems, either by treating the fairness condition as a constraint (Dwork et al., 2012) or directly optimizing the fairness metric as an objective function (Zemel et al., 2013), achieving counterfactual definitions of fairness often require an approximation of the causal model or the counterfactuals since the counterfactuals are unobservable. For example, in the FairLearning algorithm proposed by (Kusner et al., 2017), the unobserved parts of the graphical causal model are sampled through the Markov chain Monte Carlo method. Then they use only the non-descendants of sensitive variable to make the decision. However, this approach may sacrifice a significant amount of information in the data and lead to lower prediction accuracy. Moreover, when estimating the counterfactual distribution is not directly feasible due to sparse or continuous sensitive variables, unexpected relationships between sensitive and non-sensitive attributes may persist. This can either compromise fairness by considering variables related with sensitive variables, or compromise the accuracy by removing all related information regarding sensitive variables from the training data. As we discuss later, our aim is to minimize data changes while ensuring counterfactual fairness under certain assumptions, thereby better addressing the trade-off between accuracy and fairness.
Fair learning approaches
Fairness learning in machine learning aims to prevent discrimination and can be generally categorized into three stages. Firstly, pre-processing approaches (Haoyu Chen and Ghosh, 2023; Calders et al., 2009; Creager et al., 2019), which are most closely related to our work, involve modifying the data to eliminate or neutralize any preexisting bias, followed by the application of standard ML techniques. Secondly, in-processing approaches (Aghaei et al., 2019; Patil and Purcell, 2022) either by a fairness regularizer to the loss function objective, which penalizes discrimination, or imposing a constraint, thereby mitigating disparate treatment. Such methods are normally model-spefiic. The third type of approach is post-processing approaches adjust predictors learned using standard ML techniques after the fact to enhance their fairness properties (Berk et al., 2017; d’Alessandro et al., 2017; Fish et al., 2016; Hardt et al., 2016).
@@
Our approach is most related to the work by (Haoyu Chen and Ghosh, 2023), where the authors propose two distribution adjustment procedures, Orthogonalization and Marginal Distribution Mapping, for making counterfactually fair decisions based on adjusted data. While both procedures remove attributes’ dependence on sensitive variables under respective conditions, their methods provide no guarantee regarding the scale of modification to the data distribution. In contrast, our work introduces an exact approach to solving an optimization problem that guarantees minimal modification to the data while ensuring counterfactual fairness under specific assumptions. We believe that our emphasis on minimal data modification places our proposed algorithm in a unique position in the widely observed fairness-accuracy spectrum (Agarwal et al., 2018; Baharlouei et al., 2019; Berk et al., 2017). Through our approach, we aim to capture as much information as possible between the target variable $Y$ and features, including the sensitive features. Additionally, the method proposed by (Haoyu Chen and Ghosh, 2023) presuppose sensitive variables to be binary (or categorical) so that they can be easily isolated or adjusted based on empirical probability mass function. It does not address situations involving multivariate or continuous sensitive variable with intricate inter-dependency, whereas our proposed OB algorithm aims to remove the influence of a group of continuous sensitive variables from the data, thereby ensuring counterfactual fairness in subsequent machine learning tasks.
3. Methodology
3.1. Problem setup
@@
We jointly define $q$ non-sensitive variables as $A\in\mathcal{A}\subseteq\mathbb{R}^{q}$, $p$ sensitive variables as $B\in\mathbb{R}^{p}$, and decision variable as $Y\in\mathcal{Y}$. The data generation process in our problem setup can be described by a Structural Causal Model (SCM) (Pearl, 2009) as shown in Fig. 2 (a). To be specific, we consider the set of endogenous variables $V=\{B,A,Y,\hat{Y}\}$, where $\{B,A,Y\}$ are the observed variables and $\hat{Y}$ is the prediction of $Y$ we made based on $B$ and $A$. We assume that $U_{B}$, $U_{A}$, and $U_{Y}$, which are the exogenous variables that affect $B$, $A$, and $Y$ respectively, are independent of each other. The structural equations are described with the functions $F=\{f_{Y},f_{A},f_{B}\}$, one for each component in $V$, detailed as follows:
(1) $\begin{split}B&=f_{B}\left(U_{B}\right),\\ A&=f_{A}\left(B,U_{A}\right),\\ Y&=f_{Y}\left(A,U_{Y}\right).\end{split}$
@@
According to the above SCM, the bias present in the sensitive variables $B$ can transmit to the predictor $\hat{Y}$ via the non-sensitive variables $A$. This means that, if there are any differences in the distribution of $A$ conditioning on $B$, the decision variable $\hat{Y}$ based on $A$ might be unfair.
@@
In this paper, we aim to design a predictor $\hat{Y}$ that achieves the counterfactual fairness (Haoyu Chen and Ghosh, 2023; Kusner et al., 2017) without being influenced by the bias in $B$. Formally, the counterfactual fairness in our SCM can be defined as follows:
@@
Given a new pair of attributes $(\mathbf{b},\mathbf{a})$, a decision variable $Y$ is considered counterfactually fair if, for any $\mathbf{b}^{\prime}\in\mathcal{B}$,
(2) ${Y}_{\mathbf{b}^{\prime}}(U)|\left\{B=\mathbf{b}^{*},A=\mathbf{a}^{*}\right\}\stackrel{{\scriptstyle d}}{{=}}{Y}_{\mathbf{b}^{*}}(U)|\left\{B=\mathbf{b}^{*},A=\mathbf{a}^{*}\right\},$
@@
The above definition implies that the distribution of the counterfactual result should not depend on the sensitive variables conditional on the observed data. Note that although Definition 3.1 uses the decision variable $Y$, it also applies to its predictor $\hat{Y}$ without any loss of generality (Haoyu Chen and Ghosh, 2023).
3.2. Achieving counterfactual fairness via data decorrelation
To clarify and streamline the presentation of our findings, we begin by illustrating that counterfactual fairness can be attained under conditions where sensitive and non-sensitive variables exhibit no correlation and are together normally distributed.

### v2

@@ Abstract
Machine learning models have shown exceptional prowess in solving complex issues across various domains. However, these models can sometimes exhibit biased decision-making, resulting in unequal treatment of different groups. Despite substantial research on counterfactual fairness, methods to reduce the impact of multivariate and continuous sensitive variables on decision-making outcomes are still underdeveloped. We propose a novel data pre-processing algorithm, Orthogonal to Bias (OB), which is designed to eliminate the influence of a group of continuous sensitive variables, thus promoting counterfactual fairness in machine learning applications. Our approach, based on the assumption of a jointly normal distribution within a structural causal model (SCM), demonstrates that counterfactual fairness can be achieved by ensuring the data is orthogonal to the observed sensitive variables. The OB algorithm is model-agnostic, making it applicable to a wide range of machine learning models and tasks. Additionally, it includes a sparse variant to improve numerical stability through regularization. Empirical evaluations on both simulated and real-world datasets, encompassing settings with both discrete and continuous sensitive variables, show that our methodology effectively promotes fairer outcomes without compromising accuracy.
1 Introduction
@@
The fairness concern in machine learning has catalyzed a growing body of research aimed at identifying, understanding, and mitigating the biases present in data and algorithms. Among the various conceptual frameworks developed to address this issue [1, 2, 3, 4, 5, 6], the notion of counterfactual fairness [7] emerges as particularly significant. It seeks to ensure that a decision for an individual remains consistent with the decision that would have been made in a counterfactual scenario in which the individual’s sensitive variables were different. This concept is particularly powerful as it aligns closely with intuitive notions of individual fairness and justice, offering a rigorous standard by which to measure and rectify bias [8].
Significant progress has been made in achieving counterfactual fairness, with extensive studies exploring various approaches [2, 7, 9, 10, 11]. However, existing methods face two significant challenges when applied in practical scenarios: ($i$) These methods typically address discrete or binary sensitive variables and struggle to accommodate continuous variables such as age or weight [12]. ($ii$) While these methods can effectively manage a single sensitive variable, they do not extend well to scenarios involving multiple sensitive variables, thus limiting their practical utility. For instance, socio-economic status in employment hiring processes encompasses diverse factors like education, income, occupation, neighborhood, and parental education, which cannot be neatly categorized into simple discrete groups, as shown in Figure 1. Training models on historical hiring data directly may inadvertently bias them towards candidates with specific types of experience—factors often correlated with higher socio-economic status [13]. This bias may stem from unrepresentative training data or traditional fairness methods that fail to capture the complex influence of continuous multivariate sensitive variables on decision-making.
@@
Figure 1: This motivating example illustrates the distinction between binary and continuous sensitive variables in the context of hiring decisions. The dashed lines indicate the predicted hiring decisions, and the shaded indicates the fair decisions. The left panel presents a simpler scenario with a binary sensitive variable, such as gender, where adjustments for fairness are more straightforward due to limited, exhaustively enumerable values. In contrast, the right panel delves into the complexity introduced by a continuous sensitive variable, such as family income. In this case, counterfactual estimation becomes significantly more challenging due to potentially unseen or sparsely distributed values. Additionally, socio-economic status, as a multi-dimensional variable combining various factors, adds complexity to establishing fair decision-making processes.
@@
To tackle these challenges, we develop a novel data pre-processing approach that aims to remove the influence of a group of continuous sensitive variables from the data, thereby ensuring counterfactual fairness in subsequent machine learning tasks. We first prove that counterfactual fairness can be attained by making the data uncorrelated with the group of sensitive variables, based on the assumption of a jointly normal distribution within a structural causal model (SCM) framework [14]. Motivated by this understanding, we consider all sensitive variables collectively and propose a data pre-processing algorithm, referred to as Orthogonal to Bias (OB). This algorithm is designed for minimal data adjustments to achieve orthogonality between non-sensitive and sensitive data. To facilitate numerical stability, we also present a sparse variant of this algorithm to handle high-dimensional features. Then the resulting data readily serves as input for machine learning models in downstream tasks without being influenced by the undesirable bias associated with the complexities of sensitive variables. It is also important to note that our proposed algorithm is model-agnostic, making it suitable for a variety of machine learning models and tasks. Lastly, we evaluate our algorithm’s performance on two simulated datasets and three real-world datasets, covering both continuous and discrete sensitive variable settings, demonstrating that our approach enables machine learning models to achieve fairer outcomes with comparable accuracy to current state-of-the-art fair learning methods. The numerical results show that our approach robust with the joint normality assumption, indicating its applicability to a wider range of scenarios.
@@
    We show that achieving counterfactual fairness is feasible by ensuring orthogonality between non-sensitive and sensitive data under some mild conditions;
  2. 2.
@@
    We introduce a model-agnostic data pre-processing algorithm, termed Orthogonal to Bias (OB), which facilitates counterfactual fairness across a broad spectrum of downstream machine learning applications with continuous multivariate sensitive variables;
  3. 3.
@@
    We validate the enhanced efficacy of our algorithm compared to the existing state-of-the-art methods through evaluations on both synthetic and real datasets.
(a) The SCM and typical fair learning methods
(b) Transforming data with Orthogonal to Bias (OB)
@@
Figure 2: Illustration of (a) the structural causal model (SCM) and a common fair learning strategies, as well as (b) the proposed data pre-processing algorithm Orthogonal to Bias (OB). The white nodes $A$, $Y$, and $\hat{Y}$ are the non-sensitive variables, the decision variable, and its prediction, respectively. The red nodes $B$ represent the sensitive variable. The $\hat{A}$ is the transformed data that is orthogonal to bias in $B$. The gray nodes represent exogenous variables.
Related work
This section begins by examining the fundamental fairness definitions that form the basis of our modeling framework. Following this, we delve into prominent machine learning techniques designed to ensure fairness, with a particular emphasis on counterfactual fairness.
@@
Fairness in machine learning. The pursuit of fair decision-making in machine learning has led to diverse approaches for defining and quantifying fairness. Researchers commonly adopt either observational or counterfactual approaches to formalize fairness. Observational methods typically characterize fairness through metrics derived from observed data and predicted outcomes [15, 16, 10, 17]. Metrics such as individual fairness (IF) [3], demographic parity or group fairness [5, 18] and equalized odds [2] fall under this category. The key idea for the observational fairness metric is viewing fairness as treating similar individuals or individuals belonging to the same groups similarly. For example, equalized odds ensures fairness by requiring that the true positive and false positive rates are equal across different groups [9, 2].
Counterfactual methods, on the other hand, offer a causal perspective on fairness. These approaches determine fairness based on potential changes in outcomes if sensitive variables were altered [2, 7, 9, 10, 11]. Such methods extend the concept of fairness beyond observable measures. For example, the Equal Opportunity (EO) criterion directly compares the actual and counterfactual decisions for the same individual, assuming the individual had the same non-sensitive attributes, providing a more nuanced assessment than equalized odds [9].
While integrating observational fairness into machine learning models is relatively straightforward, achieving counterfactual fairness often requires approximations of causal models or counterfactuals, which presents two major challenges [9, 19]. First, the process of counterfactual estimation often compromises the predictive accuracy of models due to the exclusion of related information [20, 21, 22]. For example, the FairLearning algorithm [7] uses a Markov chain Monte Carlo method to simulate unobserved portions of the causal model, making decisions based solely on variables that are not descendants of sensitive variables. Second, estimating counterfactual distributions for sparse or continuous sensitive variables is difficult and often violates basic causal inference assumptions, namely Positivity [19]. This underscores the ongoing challenge of achieving counterfactual fairness with continuous sensitive variables. Our method addresses these two challenges by ensuring minimal data modification to achieve orthogonality between non-sensitive and sensitive data while maintaining the predictive accuracy of machine learning models.
Counterfactually fair learning. Counterfactual fairness in machine learning is achieved when a decision for an individual would remain unchanged in a counterfactual world where the individual’s sensitive variables differ [7]. This pursuit has led to the development of diverse strategies to enhance fairness, categorized into three main types: pre-processing [10], post-processing [19], and in-processing methods [12].
@@
Our approach is most related to [10], which is also a pre-processing technique for counterfactual fairness. In [10], the authors propose two distribution adjustment procedures for making counterfactually fair decisions based on adjusted data. While both procedures remove variables’ dependence on sensitive variables under respective conditions, their methods derive provide no guarantee regarding the scale of modification due to the distribution modification. In contrast, our work introduces an exact approach to solving an optimization problem that guarantees minimal modification to the data while ensuring counterfactual fairness under specific assumptions. Furthermore, the method proposed by [10] presuppose sensitive variables to be binary (or categorical) so that they can be easily isolated or adjusted based on empirical probability mass function. It does not address situations involving multivariate or continuous sensitive variable with intricate inter-dependency, whereas our proposed OB algorithm aims to remove the influence of a group of continuous sensitive variables from the data, thereby ensuring counterfactual fairness in subsequent machine learning tasks.
Additionally, our work contrasts with [12], the only other effort to the best of our knowledge that specifically tackles counterfactual fairness with continuous sensitive variables. While we include their method as a compared baseline, their approach relies on a generic loss design that lacks explicit guarantees for counterfactual fairness. In contrast, our method offers a theoretical underpinning in the motivation and design of the method, prioritizing minimal data changes to enhance the balance between accuracy and fairness.
Overall, our emphasis on minimal data modification places our proposed algorithm in a unique position within the widely observed fairness-accuracy spectrum [20, 21, 22]. Through our approach, we aim to capture as much information as possible between the target variable $Y$ and the features, including the sensitive features.
2 Methodology
2.1 Problem setup
We jointly define $q$ non-sensitive random variables as $A\in\mathcal{A}\subseteq\mathbb{R}^{q}$, $p$ sensitive variables as $B\in\mathbb{R}^{p}$, and the decision variable as $Y\in\mathcal{Y}$. The data generation process in our problem setup is described by a Structural Causal Model (SCM) [14], as shown in Figure 2(a). Our setup allows for sensitive variables ($B$) that are both continuous and multivariate, facilitating the handling of a broader range of real-world applications involving complex sensitive variables.
@@ 2.1 Problem setup
To be specific, we consider the set of endogenous variables $V=\{B,A,Y,\hat{Y}\}$, where $\{B,A,Y\}$ are the observed variables and $\hat{Y}$ is the prediction of $Y$ made based on $B$ and $A$. We assume that $U_{B}$, $U_{A}$, and $U_{Y}$, which are the exogenous variables that affect $B$, $A$, and $Y$, respectively, are independent of each other. The structural equations are described with the functions $F=\{f_{Y},f_{A},f_{B}\}$, one for each component in $V$, detailed as follows:
$B=f_{B}\left(U_{B}\right),\quad A=f_{A}\left(B,U_{A}\right),\quad Y=f_{Y}\left(A,B,U_{Y}\right).\quad$ (1)
@@
According to the above SCM, the bias present in the sensitive variables $B$ can transmit to the predictor $\hat{Y}$ via the non-sensitive variables $A$. This means that if there are any differences in the distribution of $A$ conditioned on $B$, the decision variable $\hat{Y}$ based on $A$ might be unfair.
@@
In this paper, we aim to design a predictor $\hat{Y}$ that achieves counterfactual fairness [7, 10] without being influenced by the bias in $B$. Formally, the counterfactual fairness in our SCM can be defined as follows:
@@
Given a new pair of variables $(\mathbf{b},\mathbf{a})$, a decision variable $Y$ is considered counterfactually fair if, for any $\mathbf{b}^{\prime}\in\mathcal{B}$,
${Y}_{\mathbf{b}^{\prime}}(U)|\left\{B=\mathbf{b}^{*},A=\mathbf{a}^{*}\right\}\stackrel{{\scriptstyle d}}{{=}}{Y}_{\mathbf{b}^{*}}(U)|\left\{B=\mathbf{b}^{*},A=\mathbf{a}^{*}\right\},$ (2)
@@
The above definition implies that the distribution of the counterfactual result should not depend on the sensitive variables conditional on the observed data. Note that although Definition 2.1 uses the decision variable $Y$, it also applies to its predictor $\hat{Y}$ without any loss of generality [10].
2.2 Achieving counterfactual fairness via data decorrelation
To clarify and streamline the presentation of our findings, we begin by showing that counterfactual fairness can be attained under conditions where sensitive and non-sensitive variables exhibit no correlation and are joint normal.